# Role of $\Sigma^{*}(1385)$ on $\Lambda$ hyperon polarization in relativistic heavy ion collisions

### v1

@@ Abstract
The effect of $\Sigma^{*}(1385)$ baryon resonance on the time evolution of the $\Lambda$ hyperon polarization in hadronic matter is studied using a kinetic approach. This approach explicitly includes the production of the $\Sigma^{*}$ resonance from the $\Lambda-\pi$ and $\Sigma(1192)-\pi$ scatterings as well as its decay into $\Lambda+\pi$ or $\Sigma+\pi$. The resulting coupled kinetic equations governing the time evolution of $\Lambda$, $\Sigma$ and $\Sigma^{*}$ numbers and polarizations are solved for Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% centrality, using initial values determined by thermal yields and the thermal vorticity at chemical freeze-out temperature. As the hadronic matter expands and cools, the $\Lambda$ polarization is found to increase slightly during early times and then decreases very slowly afterwards, while the $\Sigma$ polarization remains nearly constant and the $\Sigma^{*}$ polarization continuously decreases. Including feed-down contributions to the $\Lambda$ polarization from the decays of partially polarized $\Sigma^{0}$, $\Sigma^{*}$, and $\Xi(1322)$ hyperons, where the $\Xi$ polarization is obtained by solving coupled kinetic equations for the $\Xi$ and $\Xi^{*}(1532)$ system, the resulting $\Lambda$ polarization becomes smaller and decreases over time. In both cases, however, the time variation of the $\Lambda$ polarization is sufficiently small to support the assumption of an early freeze-out of $\Lambda$ spin degree of freedom in relativistic heavy ion collisions.
@@ I Introduction
Since the STAR Collaboration’s 2017 observation that the $\Lambda$ hyperon produced in heavy ion collisions at the Relativistic Heavy Ion Collider (RHIC) are partially polarized along the direction perpendicular to the reaction plane Adamczyk et al. (2017a), numerous theoretical studies have been carried out to understand the origin of this phenomenon Becattini and Lisa (2020). Most of these studies are based on hydrodynamic models Karpenko and Becattini (2017); Fu et al. (2021); Becattini et al. (2021a); Palermo et al. (2024) or the course-grained transport model Li et al. (2017), which assume thermal equilibrium between the $\Lambda$ spin and the vortical fluid. In these approaches, the global $\Lambda$ polarization is calculated from the thermal vorticity at the chemical freeze-out temperature.
Extensions of these studies have incorporated the effects of thermal shear in the fluid to address the measured $\Lambda$ local polarization Niida (2019), particularly its component along the beam direction, as a function of the azimuthal angle in the transverse plane of the collision Liu and Yin (2021); Becattini et al. (2021b); Yi et al. (2021).
Beyond equilibrium models, the $\Lambda$ polarization has been studied using non-equilibrium framework such as the chiral kinetic approach Sun and Ko (2017a, 2019) and the covariant angular-momentum-conserved chiral transport model Liu et al. (2020). In these approaches, both vorticity and shear in the quark matter affect the dynamics of quarks and antiquarks through their equations of motion and scatterings. The polarization of the $\Lambda$ hyperon is then determined via the coalescence of polarized strange quarks with polarized up and down quarks during hadronization of the quark matter Sun and Ko (2017b).
@@
[…]
However, both of these studies are based on the assumption of instantaneous $\Lambda-\pi$ resonance scattering via the $\Sigma^{*}$ resonance, and neglect feed-down contributions to the $\Lambda$ polarization from other hadrons. Since the experimentally measured $\Lambda$ polarization includes not only primary $\Lambda$ hyperons but also those from the decays of resonances Becattini et al. (2017); Xia et al. (2019) – particularly the strong decay $\Sigma^{*}\to\Lambda+\pi$, the electromagnetic decay $\Sigma^{0}(1192)\to\Lambda+\gamma$, and the weak decays $\Xi^{0}(1315)\to\Lambda+\pi^{0}$ and $\Xi^{-}(1322)\to\Lambda+\pi^{-}$ – these contributions must be considered.
In the present study, we go beyond the the approximation used in the kinetic approach of Ref. Sung et al. (2024) by treating the $\Sigma^{*}$ resonance dynamically via the processes $\Lambda+\pi\leftrightarrow\Sigma^{*}$. This includes the production of $\Sigma^{*}$ from $\Lambda-\pi$ scattering and its decay back to $\Lambda$ and $\pi$, as well as its decay to $\Sigma+\pi$ and its inverse $\Sigma+\pi\to\Sigma^{*}$ process.
In addition, we consider feed-down contributions to the $\Lambda$ polarization from the decays of $\Sigma^{0}$ and $\Sigma^{*}$ as well as from $\Xi^{0}$ and $\Xi^{-}$, by solving coupled kinetic equations for the $\Xi$ and $\Xi^{*}(1532)$ hyperons during the hadronic stage of relativistic heavy ion collisions.
@@ II $\Lambda+\pi\leftrightarrow\Sigma^{*}\leftrightarrow\Sigma+\pi$ processes
×$\hskip 34.14322pt\times\frac{B_{\Sigma^{*}\to Y+\pi}s\Gamma_{\Sigma^{*}}^{2}(s)}{(s-m_{\Sigma^{*}}^{2})^{2}+s\Gamma_{\Sigma^{*}}^{2}(s)},\quad Y=\Lambda,\Sigma.$ (1)
@@
×$\hskip 14.45377pt\times\int_{z_{0}}^{\infty}dzK_{1}(z)\sigma_{Y+\pi\to\Sigma^{*}}(z^{2}T^{2})$
×$\hskip 14.45377pt\times[z^{2}-(m_{Y}+m_{\pi})^{2}/T^{2}][z^{2}-(m_{Y}-m_{\pi})^{2}/T^{2}].$ (3)
In the above, $f_{Y}({\bf p}_{Y})$ is the Boltzmann momentum distribution, given by $f_{Y}({\bf p}_{Y})=e^{-(\sqrt{m_{Y}^{2}+{\bf p}_{Y}^{2}}-\mu_{Y})/T}$, where the chemical potential $\mu_{Y}=B_{Y}\mu_{B}+S_{Y}\mu_{S}$ is expressed in terms of the baryon number $B_{Y}$ and the strangeness number $S_{Y}$ of the hyperon, along with the baryon chemical potential $\mu_{B}$ and strangeness chemical potential $\mu_{S}$ of the hadronic matter.
The relative velocity between $Y$ and $\pi$ in Eq.(III) is given by $v=\sqrt{(p_{Y}\cdot p_{\pi})^{2}-m_{Y}^{2}m_{\pi}^{2}}/(E_{Y}E_{\pi})$, where $p_{Y,\pi}=(E_{Y,\pi},{\bf p}_{Y,\pi})$ are their four momenta. The functions $K_{1}(z)$ and $K_{2}(m_{i}/T)$ are the modified Bessel functions of the first and second kinds, respectively.
@@ IV Kinetic equations for $\Lambda$, $\Sigma$, and $\Sigma^{*}$ numbers and polarizations
by taking appropriate combinations of the kinetic equations in Eq.(IV). The resulting coupled equations are given by,
@@
×$\hskip 36.98866pt\times\left(\frac{5}{9}P_{Y}-P_{\Sigma^{*}}\right).$ (8)
Since the right-hand side of the kinetic equations in Eq.(IV) vanish when hyperons reach their thermally equilibrated values, given by $z_{Y}^{(T)}N_{Y}^{(T)}$, where $z_{Y}^{(T)}$ represents the hyperon fugacity. In this case, terms involving $\langle\Gamma_{\Sigma^{*}\to Y+\pi}\rangle$ can be related to $\langle\sigma_{Y+\pi\to\Sigma^{*}}v\rangle$ through detailed balance, i.e.,
@@
[…]
×$\hskip 34.14322pt\times\frac{g_{\Sigma^{*}}V}{(2\pi)^{3}}\int d^{3}{\bf p}e^{-(\sqrt{s+{\bf p}^{2}}-\mu_{\Sigma^{*}})/T},$ (10)
@@
×$\hskip 36.98866pt\times\left(N_{Y}-\frac{N_{Y}^{(T)}}{z_{\pi}^{(T)}N_{\Sigma^{*}}^{(T)}}N_{\Sigma^{*}}\right),$
@@
×$\hskip 36.98866pt\times\frac{N_{Y}}{N_{\Sigma^{*}}}\left(\frac{5}{9}P_{Y}-P_{\Sigma^{*}}\right).$ (11)
@@
[…]
@@
Figure 3: Time evolution of $\Lambda$ (solid line), $\Sigma$ (dashed line), and $\Sigma^{*}(1385)$ (dotted line) polarizations in Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% collision centrality, for initial $\Lambda$ and $\Sigma$ polarizations of 0.025 and initial $\Sigma^{*}$ polarization of 0.042. Also shown (dash-dotted line) is the $\Lambda$ polarization $P_{\Lambda}^{\rm FD}$ that includes the feed-down contributions described in the text.
@@
The time evolution of $\Lambda$, $\Sigma$, and $\Sigma^{*}$ polarizations, obtained by solving the last two equations of Eq.(IV) with the time-dependent $\Lambda$, $\Sigma$ and $\Sigma^{*}$ numbers from Fig. 2, is shown in Fig. 3. The $\Lambda$ polarization increases slightly with time and reaches about 0.028 at $T_{K}$ when the $\Sigma^{*}$ polarization decreases from its initial value of 0.042. Compared with the $\sim$7decrease in $\Lambda$ polarization found in Ref. Sung et al. (2024) when the $\Sigma^{*}$ resonance was not treated as a dynamic degree of freedom, the present results show a slightly enhancement of $\Lambda$ polarization due to hadronic scattering effects during the hadronic stage of relativistic heavy ion collisions. If the hadronic stage were to last twice as long, ending at $\tau=8~{\rm fm}/c$ when the temperature has dropped to $T=112$ MeV, the $\Lambda$ polarization remains to have essentially the same value of 0.028. The $\Sigma^{*}$ polarization, on the other hand, continuously decreases with time to the value 0.024. In contrast, the time evolution of the $\Sigma$ polarization is minimal due to the smaller $\Sigma+\pi\to\Sigma^{*}$ cross section compared to that of $\Lambda+\pi\to\Sigma^{*}$.
@@ VI Feed-down contributions from $\Xi$ and $\Xi^{*}$ to $\Lambda$ polarization
×$\hskip 36.98866pt\times\left(N_{\Xi}-\frac{N_{\Xi}^{(T)}}{z_{\pi}^{(T)}N_{\Xi^{*}}^{(T)}}N_{\Xi^{*}}\right),$
@@
Figure 6: Time evolution of $\Xi$ (solid line) and $\Xi^{*}$ (dotted line) polarizations in Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% collision centrality, with initial $\Xi$ polarizations of 0.025 and $\Xi^{*}$ polarization of 0.042.
@@
where the $C_{Y}$ are the polarization transfer coefficients of hyperons to the $\Lambda$ in their decay. Their values for $\Lambda$, $\Sigma^{0}$, $\Xi^{0}$, and $\Xi^{-}$ are $C_{\Lambda}=1$, $C_{\Sigma^{0}}=-1/3$, $C_{\Xi^{0}}=0.900$, and $C_{\Xi^{-}}=0.927$ Becattini et al. (2017); Xia et al. (2019). For $\Sigma^{{}^{+}}$, $\Sigma^{*0}$, and $\Sigma^{*-}$, their polarization transfer coefficients are determined as follows. The $\Sigma^{*0}$ has the strong decay channels $\Sigma^{*0}\to\Lambda+\pi^{0}$, $\Sigma^{*0}\to\Sigma^{+}+\pi^{-}$, and $\Sigma^{*0}\to\Sigma^{-}+\pi^{+}$ but only the first one contributes to the $\Lambda$ polarization, with $C_{\Sigma^{*0}}=0.88$. For $\Sigma^{*\pm}$, the strong decays include $\Sigma^{*\pm}\to\Lambda+\pi^{\pm}$, $\Sigma^{*\pm}\to\Sigma^{\pm}+\pi^{0}$, and $\Sigma^{*\pm}\to\Sigma^{0}+\pi^{\pm}$. Taking into account the branching ratios of 88% and 12% for $\Sigma^{*}$ decaying to $\Lambda$ and $\Sigma$, respectively, and the equal decay probabilities to charged and neutral $\Sigma$, we obtain the polarization transfer coefficients $C_{\Sigma^{*+}}=C_{\Sigma^{*-}}=0.88+0.12\times\frac{1}{2}\times(-\frac{1}{3})=0.86$, since only $\Sigma^{0}$ can decay electromagnetically to $\Lambda+\gamma$, with a polarization transfer coefficient of 1/3-1/3- 1 / 3.
[…]

### v2

# Role of $\Sigma^{*}(1385)$ in $\Lambda$ hyperon polarization in relativistic heavy ion collisions

@@ Abstract
The effect of $\Sigma^{*}(1385)$ baryon resonance on the time evolution of the $\Lambda$ hyperon polarization in hadronic matter is studied using a spin-dependent kinetic approach. This approach explicitly includes the production of the $\Sigma^{*}$ resonance from the $\Lambda-\pi$ and $\Sigma(1192)-\pi$ scatterings as well as its decay into $\Lambda+\pi$ or $\Sigma+\pi$, all including explicitly their spin dependence. The resulting coupled kinetic equations governing the time evolution of $\Lambda$, $\Sigma$ and $\Sigma^{*}$ numbers and polarizations are solved for Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% centrality, using initial values determined by their thermal yields and the thermal vorticity in the hadronic matter at chemical freeze-out temperature. As the hadronic matter expands and cools, the $\Lambda$ polarization is found to increase slightly during early times and then decreases very slowly afterwards, while the $\Sigma$ polarization remains nearly constant and the $\Sigma^{*}$ polarization continuously decreases. Including feed-down contributions to the $\Lambda$ polarization from the decays of partially polarized $\Sigma^{0}$, $\Sigma^{*}$, and $\Xi(1322)$ hyperons, where the $\Xi$ polarization is obtained by solving coupled kinetic equations for the $\Xi$ and $\Xi^{*}(1532)$ system, the resulting $\Lambda$ polarization becomes smaller and decreases over time. In both cases, the time variation of the $\Lambda$ polarization is, however, sufficiently small to support the assumption of an early freeze-out of $\Lambda$ spin degree of freedom in relativistic heavy ion collisions.
@@ I Introduction
Since the observation by the STAR Collaboration in 2017 that the $\Lambda$ hyperon produced in heavy ion collisions at the Relativistic Heavy Ion Collider (RHIC) are partially polarized along the direction perpendicular to the reaction plane Adamczyk et al. (2017a), numerous theoretical studies have been carried out to understand the origin of this phenomenon Becattini and Lisa (2020). Most of these studies are based on hydrodynamic models Karpenko and Becattini (2017); Fu et al. (2021); Becattini et al. (2021a); Palermo et al. (2024) or the course-grained transport model Li et al. (2017), which assume thermal equilibrium between the $\Lambda$ spin and the vortical fluid, and calculate the global $\Lambda$ polarization from the thermal vorticity at the chemical freeze-out temperature. Extensions of these studies have incorporated the effects of thermal shear in the fluid to address the measured $\Lambda$ local polarizations Niida (2019), particularly its component along the beam direction, as functions of the azimuthal angle in the transverse plane of the collision Liu and Yin (2021); Becattini et al. (2021b); Yi et al. (2021).
Besides equilibrium models, the $\Lambda$ polarization has also been studied using non-equilibrium framework such as the chiral kinetic approach Sun and Ko (2017a, 2019) and the covariant angular-momentum-conserved chiral transport model Liu et al. (2020). In these studies, both vorticity and shear in the quark matter affect the dynamics of quarks and antiquarks through their equations of motion and scatterings. The polarization of the $\Lambda$ hyperon is then determined via the coalescence of polarized strange quarks with polarized up and down quarks during hadronization of the quark matter Sun and Ko (2017b).
@@
As proposed in Ref. Ko (2023), this early freeze-out would occur if $\Lambda-\pi$ scattering proceeds through the spin 3/2, positive parity $\Sigma^{*}(1385)$ resonance, since the ratio of $\Lambda$ spin non-flip to flip probabilities in such scattering is 3.5. This explanation has been supported by detailed theoretical calculations of the $\Lambda$ polarization relaxation time in hot hadronic matter, which is found to be about 4-7 fm/$c$  Hidaka et al. (2024), and by the kinetic approach to the time evolution of $\Lambda$ polarization during the hadronic stage of relativistic heavy ion collisions, which show a decreases of about 7-12% Sung et al. (2024). However, these studies are based on the approximation of instantaneous $\Lambda-\pi$ resonance scattering via the $\Sigma^{*}$ resonance, and neglect feed-down contributions to the $\Lambda$ polarization from other hadrons. Since the experimentally measured $\Lambda$ polarization includes not only primary $\Lambda$ hyperons but also those from the decays of resonances Becattini et al. (2017); Xia et al. (2019) – particularly the strong decay $\Sigma^{*}\to\Lambda+\pi$, the electromagnetic decay $\Sigma^{0}(1192)\to\Lambda+\gamma$, and the weak decays $\Xi^{0}(1315)\to\Lambda+\pi^{0}$ and $\Xi^{-}(1322)\to\Lambda+\pi^{-}$ – these contributions must be considered.
In the present study, we go beyond the approximation used in the kinetic approach of Ref. Sung et al. (2024) by treating the $\Sigma^{*}$ resonance dynamically via the processes $\Lambda+\pi\leftrightarrow\Sigma^{*}$, i.e., the production of $\Sigma^{*}$ from $\Lambda-\pi$ scattering and its decay back to $\Lambda$ and $\pi$. In addition, we consider feed-down contributions to the $\Lambda$ polarization from the decays of $\Sigma^{0}$ and $\Sigma^{*}$ as well as from $\Xi^{0}$ and $\Xi^{-}$, with the latter obtained by solving coupled kinetic equations for the numbers and polarizations of $\Xi$ and $\Xi^{*}(1532)$ hyperons during the hadronic stage of relativistic heavy ion collisions.
@@ II $\Lambda+\pi\leftrightarrow\Sigma^{*}\leftrightarrow\Sigma+\pi$ processes
$\hskip 34.14322pt\times\frac{B_{\Sigma^{*}\to Y+\pi}s\Gamma_{\Sigma^{*}}^{2}(s)}{(s-m_{\Sigma^{*}}^{2})^{2}+s\Gamma_{\Sigma^{*}}^{2}(s)},\quad Y=\Lambda,\Sigma.$ (1)
@@
$\hskip 14.45377pt\times\int_{z_{0}}^{\infty}dzK_{1}(z)\sigma_{Y+\pi\to\Sigma^{*}}(z^{2}T^{2})$
$\hskip 14.45377pt\times[z^{2}-(m_{Y}+m_{\pi})^{2}/T^{2}][z^{2}-(m_{Y}-m_{\pi})^{2}/T^{2}].$ (3)
In the above, $f_{Y}({\bf p}_{Y})$ is the Boltzmann momentum distribution, given by $f_{Y}({\bf p}_{Y})=e^{-(\sqrt{m_{Y}^{2}+{\bf p}_{Y}^{2}}-\mu_{Y})/T}$, where the chemical potential $\mu_{Y}=B_{Y}\mu_{B}+S_{Y}\mu_{S}$ is expressed in terms of the baryon number $B_{Y}$ and the strangeness number $S_{Y}$ of the hyperon, along with the baryon chemical potential $\mu_{B}$ and strangeness chemical potential $\mu_{S}$ of the hadronic matter. The relative velocity between $Y$ and $\pi$ in Eq.(III) is given by $v=\sqrt{(p_{Y}\cdot p_{\pi})^{2}-m_{Y}^{2}m_{\pi}^{2}}/(E_{Y}E_{\pi})$, where $p_{Y,\pi}=(E_{Y,\pi},{\bf p}_{Y,\pi})$ are their four momenta. The functions $K_{1}(z)$ and $K_{2}(m_{i}/T)$ are the modified Bessel functions of the first kind.
@@ IV Kinetic equations for $\Lambda$, $\Sigma$, and $\Sigma^{*}$ numbers and polarizations
by taking appropriate combinations of the kinetic equations in Eq.(IV). The resulting coupled kinetic equations are given by,
@@
$\hskip 36.98866pt\times\left(\frac{5}{9}P_{Y}-P_{\Sigma^{*}}\right).$ (8)
Since the right-hand side of the kinetic equations in Eq.(IV) vanish when hyperons reach their thermally equilibrated values, given by $z_{Y}^{(T)}N_{Y}^{(T)}$, with $z_{Y}^{(T)}$ being the hyperon fugacity, terms involving $\langle\Gamma_{\Sigma^{*}\to Y+\pi}\rangle$ can be related to $\langle\sigma_{Y+\pi\to\Sigma^{*}}v\rangle$ through detailed balance, i.e.,
@@
$N_{\Sigma^{*}_{m}}^{(T)}=\frac{A_{\Sigma^{*}}}{\pi}\int_{s_{0}}^{\infty}ds\frac{\sqrt{s}\Gamma_{\Sigma^{*}}(\sqrt{s})}{(s-m_{\Sigma^{*}}^{2})^{2}+s\Gamma_{\Sigma^{*}}^{2}(\sqrt{s})}$
$\hskip 34.14322pt\times\frac{g_{\Sigma^{*}}V}{(2\pi)^{3}}\int d^{3}{\bf p}e^{-(\sqrt{s+{\bf p}^{2}}-\mu_{\Sigma^{*}})/T},$ (10)
@@
$\hskip 36.98866pt\times\left(N_{Y}-\frac{N_{Y}^{(T)}}{z_{\pi}^{(T)}N_{\Sigma^{*}}^{(T)}}N_{\Sigma^{*}}\right),$
@@
$\hskip 36.98866pt\times\frac{N_{Y}}{N_{\Sigma^{*}}}\left(\frac{5}{9}P_{Y}-P_{\Sigma^{*}}\right).$ (11)
@@
With the chemical and kinetic freeze-out times obtained from the AMPT model simulations of these collisions Xu and Ko (2017), namely $\tau_{C}=5.13$ fm/$c$ and $\tau_{K}=6.46$ fm/$c$, the time evolution of the temperature and volume of the hadronic matter can be parametrized as follows:
@@
Figure 3: Time evolution of $\Lambda$ (solid line), $\Sigma$ (dashed line), and $\Sigma^{*}(1385)$ (dotted line) polarizations in Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% collision centrality, for initial $\Lambda$ and $\Sigma$ polarizations of 0.025 and initial $\Sigma^{*}$ polarization of 0.042. Also shown (dot-dashed line) is the $\Lambda$ polarization $P_{\Lambda}^{\rm FD}$ that includes the feed-down contributions described in the text.
@@
The time evolution of $\Lambda$, $\Sigma$, and $\Sigma^{*}$ polarizations, obtained by solving the last two equations of Eq.(IV) with the time-dependent $\Lambda$, $\Sigma$ and $\Sigma^{*}$ numbers from Fig. 2, is shown in Fig. 3. The $\Lambda$ polarization increases slightly with time and reaches about 0.028 at $T_{K}$ when the $\Sigma^{*}$ polarization decreases from its initial value of 0.042 to 0.03 at $T_{K}$. Compared with the $\sim$7% decrease in $\Lambda$ polarization found in Ref. Sung et al. (2024) when the $\Sigma^{*}$ resonance was not treated as a dynamic degree of freedom, the present results show a slightly enhancement of $\Lambda$ polarization due to hadronic scattering effects during the hadronic stage of relativistic heavy ion collisions. If the hadronic stage were to last twice as long, ending at $\tau=8~{\rm fm}/c$ when the temperature has dropped to $T=112$ MeV, the $\Lambda$ polarization remains to have essentially the same value of 0.028. The $\Sigma^{*}$ polarization, on the other hand, continuously decreases with time to the value 0.024. In contrast, the time evolution of the $\Sigma$ polarization is minimal due to the smaller $\Sigma+\pi\to\Sigma^{*}$ cross section compared to that of $\Lambda+\pi\to\Sigma^{*}$.
@@ VI Feed-down contributions from $\Xi$ and $\Xi^{*}$ to $\Lambda$ polarization
$\hskip 36.98866pt\times\left(N_{\Xi}-\frac{N_{\Xi}^{(T)}}{z_{\pi}^{(T)}N_{\Xi^{*}}^{(T)}}N_{\Xi^{*}}\right),$
@@
Figure 6: Time evolution of $\Xi$ (solid line) and $\Xi^{*}$ (dotted line) polarizations in Au-Au collisions at $\sqrt{s_{NN}}=7.7$ GeV and 20-50% collision centrality, with initial $\Xi$ polarizations of 0.025 and $\Xi^{*}$ polarization of 0.042. The dot-dashed line denotes the $\Xi$ polarization $P_{\Xi}^{\rm FD}$ that includes the feed-down contribution from $\Xi^{*}$.
@@
where the $C_{Y}$ are the polarization transfer coefficients of hyperons to the $\Lambda$ in their decay. Their values for $\Lambda$, $\Sigma^{0}$, $\Xi^{0}$, and $\Xi^{-}$ are $C_{\Lambda}=1$, $C_{\Sigma^{0}}=-1/3$, $C_{\Xi^{0}}=0.900$, and $C_{\Xi^{-}}=0.927$ Becattini et al. (2017); Xia et al. (2019). For $\Sigma^{{}^{+}}$, $\Sigma^{*0}$, and $\Sigma^{*-}$, their polarization transfer coefficients are determined as follows. The $\Sigma^{*0}$ has the strong decay channels $\Sigma^{*0}\to\Lambda+\pi^{0}$, $\Sigma^{*0}\to\Sigma^{+}+\pi^{-}$, and $\Sigma^{*0}\to\Sigma^{-}+\pi^{+}$ but only the first one contributes to the $\Lambda$ polarization, with $C_{\Sigma^{*0}}=0.88$. For $\Sigma^{*\pm}$, the strong decays include $\Sigma^{*\pm}\to\Lambda+\pi^{\pm}$, $\Sigma^{*\pm}\to\Sigma^{\pm}+\pi^{0}$, and $\Sigma^{*\pm}\to\Sigma^{0}+\pi^{\pm}$. Taking into account the branching ratios of 88% and 12% for $\Sigma^{*}$ decaying to $\Lambda$ and $\Sigma$, respectively, and the equal decay probabilities to charged and neutral $\Sigma$, we obtain the polarization transfer coefficients $C_{\Sigma^{*+}}=C_{\Sigma^{*-}}=0.88+0.12\times\frac{1}{2}\times(-\frac{1}{3})=0.86$, since only $\Sigma^{0}$ can decay electromagnetically to $\Lambda+\gamma$, with a polarization transfer coefficient of 1/3-1/3.
Using the time dependence of $\Lambda$, $\Sigma$, and $\Sigma^{*}$ numbers in Fig. 2 and of the $\Xi$ number in Fig. 5, which give $N_{\Sigma^{0}}=N_{\Sigma}/3$ and $N_{\Xi^{0}}=N_{\Xi^{-}}=N_{\Xi}/2$, as well as the time dependence of their polarizations from Figs. 3 and 6, the resulting time dependence of $P_{\Lambda}^{\rm FD}$ is shown by the dot-dashed line in Fig. 3. It is seen that $P_{\Lambda}^{\rm FD}$ has an initial value of about 0.02 at $\tau_{C}=5.13$ fm/$c$, decreases to 0.019 at $\tau_{K}=6.46$ fm/$c$, and becomes 0.017 at $\tau=8$ fm/$c$, which are smaller than the value of $P_{\Lambda}$ by 20%, 34%, and 39% at corresponding times.
@@
The above conclusion is also true for the $\Xi$ polarization because it only increases by less than 10% during the hadronic stage as shown by the solid line in Fig. 6. Including the feed-down contribution from the $\Xi^{*}$ due to the decay $\Xi^{*}\to\Xi+\pi$ does not change this conclusion as shown by the dot-dashed line in the same figure. The latter, denoted by $P_{\Xi}^{\rm FD}$, is calculated according to
$P_{\Xi}^{\rm FD}=\frac{P_{\Xi}N_{\Xi}+P_{\Xi^{*}}N_{\Xi^{*}}}{N_{\Xi}+N_{\Xi^{*}}},$ (15)
[…]
Our spin-dependent kinetic approach can further be used to study the hadronic effects on the polarization of $\Omega(1672)$ hyperons, e.g. by considering their scatterings with pions and kaons through its resonances Navas et al. (2024). Such a study will allow us to check if the $\Omega(1672)$ polarization also freezes out earlier as assumed in Ref. Wei et al. (2019), which is essential for making meaningful comparisons with the measurements by the STAR Collaboration Adam et al. (2021).